OpenStudy (anonymous):

for an art project you make a square print with side length of 8 inches. you make a frame using stips of wood 1 1/4 inches wide. what is the area of the frame

OpenStudy (anonymous):

Let's consider the drawing  Normally, the area of a square is \(s^2\) where \(s\) is the length of a side. So in this case, we need to calculate s. We know \(s\geq8\) Because it is initially 8 inches. So the inner square area is \(8^2units\) The outer square area is: \((8+2(\frac{5}{4}))^2\) So therefore, the area of just the frame is: \[\left[8+\left(2\times\frac{5}{4}\right)\right]^2-8^2\]

OpenStudy (anonymous):

I will leave the calculation to you